Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 244     Entry time: 11 Jun 2023, 02:08
Author: Ulrike Boesenberg 
Group: MID 
Subject: Shift summary 

* First half of the day was used to optimize seeding performance
* Cryo issues started at 6 pm; were resolved by 9:30pm
* After some tuning, we observed very nice seeding performance around midnight (ratio better than 1:1)
* Beam size seems too big; we measured 25x25 um w/ knife edge scans and CRL2 lenses 3,5,6
* Mounted the x-ray again, and left the chamber pumping overnight
* Morning crew, please double-check the beam size tomorrow

ELOG V3.1.4-7c3fd00